In a hydrogen atom, if energy of an electron in ground state is -13.6ev , then that in 2nd excited state is:(a) -1.51eV(b) -3.4eV(c) -6.04eV(d) -13.6eV

Grade:11

2 Answers

Khushi goyal
6 years ago
E=-13.6*z^2/n^2eV Where z = atomic numberAnd n=no. Of orbitEnergy in 2nd excited state =-13.6×1^2/2^2=-13.6eV
Yeshwanth Lakshman
6 years ago
2nd excited state means that n=3En= -13.6z²eV/n² = -13.6ev/n² (z=1 for hydrogen) = -13.6ev/9 = -1.51ev